OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 12x - y = -4 4x - 3y = -6 4x - 3(-12x - 4) = -6 4x - 3(12x + 4) = -6 4(12x + 4) - 3y = -6 4(-12x - 4) - 3y = -6

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

Do you know how to isolate y by itself on one side of the equation in the first equation?

OpenStudy (anonymous):

4x - 3(12x + 4) = -6

OpenStudy (anonymous):

12x - y = -4 -12x -12x -y = -4 - 12x divide by -1 y = 4 + 12x @Cosmichaotic

OpenStudy (anonymous):

Perfect! Ok give me one sec here.

OpenStudy (anonymous):

i already got the answer but thank you

OpenStudy (anonymous):

Ok so now that we know what "y" is equal too. (You found 'y' to be equal to 4 + 12x or y = 4 + 12x) We plug in 4 + 12x for y in the 2nd equation. 4x - 3y = -6 (Second Equation) Can you show me what it would look like after you do this?

OpenStudy (anonymous):

Tip: When you substitute like this, try to always use parentheses around what you are plugging in. y = 4 + 12x When you plug it in for y, make sure you plug in (4 + 12x) with parentheses. It will help you do the math correctly and more easily.

OpenStudy (anonymous):

Go ahead and plug in 4 + 12x for y into 4x - 3y = -6

OpenStudy (anonymous):

Oh alright, well would you like me to continue so you can see how to do this in the future for tests, quizzes, and more importantly for your own confidence? =0)

OpenStudy (anonymous):

If not, that's alright.

OpenStudy (anonymous):